Find the general solution of y′′ + 2y = sin(7t/5)

Please show all steps and work clearly so I can follow your logic and learn to solve similar ones myself. Will rate your answer. Thank you kindly!
Answer #1

y + 2y = sm / h (D++2) y = Sum(75) r²+2=0 = - r²= -2 of-+iva y = q (as (2 t) + (2 sin(Jet) (1) سن و - 4 - t2 men (3) 25 sin/